What are the coordinates of the vertex for y=-3x^2+24x+3
Thank you!

Answer by stanbon(75887)

What are the coordinates of the vertex for y=-3x^2+24x+3
------
Vertex occurs where x = -b/(2a) = -24/(2*-3) = 4
-----
f(4) = -3*16 + 24*4 + 3 = -48 + 96 + 3 = 48 + 3 = 51
------------------
Vertex:: (4,51)
------------
Cheers,
Stan H.
-----------